OpenStudy (anonymous):

can anyone help me solve systems of equations with 3 varibles... the first equation is 2x + 2y +3z=14 second equation is 3x + y - z=-3 third equation x + y +2z =9 any help greatly appreciated

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Can be done using Cramers rule or the determinant method

OpenStudy (amistre64):

x + y +2z = 9 [*-2] x + y +2z = 9 [*-3] 2x +2y +3z =14 3x + y - z = -3 -2x -2y -4z = -18 -3x -3y -6z = -27 2x +2y +3z = 14 3x + y - z = -3 ----------------- ------------------ -z = -4 -2y -7z = -30 z = 4 -2y -7(4) = -30 -2y -28 = -30 -2y = -2 y=1

OpenStudy (amistre64):

x + y +2z = 9 x + 1 +2(4) = 9 x = 0 dbl chk

OpenStudy (amistre64):

i think it checks .... but i tend to be biased :)